OpenStudy (anonymous):

MEDAL FOR BEST ANSWER A six-sided number cube is rolled. Find the probability of rolling a 3 or a number greater than 4. A. 1/18 B. 2/3 C. 1/6 D. 1/2

OpenStudy (anonymous):

Ok. So you roll a number cube with 6 different possible outcomes. 1 2 3 4 5 6 We want to know what the probability is of rolling a prime number or a number greater than 4. To do this, find the probability of each, and then add them.

OpenStudy (anonymous):

Okay, so the prime numbers between 1 and 6 are 2,3, and 5. So there are 3 prime numbers. The chance of rolling a prime number is 3 out of 6, or 1/2. And the numbers greater than 4 are 5 and 6 (4 isn't greater than 4). So there are 2 numbers greater than 4. The chance of rolling a number greater than 4 is 2 out of 6, or 1/3.

OpenStudy (anonymous):

So add 1/2 and 1/3. 1/2+1/3 3/6+2/6=5/6. So the chance of rolling a prime number or a number greater than 4 is 5/6. But each probability has 5 as a possibility, so subtract 1/6. 5/6-1/6=4/6 or 2/3.

OpenStudy (anonymous):

Thank you very much!!

OpenStudy (anonymous):

np hmu for more help